# Optimizing Power Grid Topologies with Reinforcement Learning: A Survey of Methods and Challenges

### v1

@@ 3.4 Observation space
Active power Reactive power Voltages Cooldown Solution Acronym Complete Observation Time stamp Loads ($load\_p$) Gens ($gen\_p$) Lines ($p\_ex,p\_or$) Loads ($load\_qp$) Gens ($gen\_q$) Lines ($q\_ex,q\_or$) Loads ($load\_v$) Gens ($gen\_v$) Lines ($v\_ex,v\_or$) Line flows ($a\_{or},a\_{ex}$) Line loads ($\rho$) Time step overflow Line status Topo config ($topo\_vect$) line sub Maintenance Extra A3C-2019 DDQN-2019 A3C-2020181818Based on code, not (clearly) documented SMAAC-202118 Danger and history D3QN-202018 Binbinchen-2020 SAS-202118 CEM-2021 D3QN-2022 PowRL-2022 AlphaZero-2022 Curriculum-202318 HRL-2023 MARL-2023 Danger and history LJNAgent-202418 IL-2024 Thermal limit BDQN-2024 Zonal-2024 GNNIL-2025 Thermal limit CCMA-2025 Actions and action values.
@@ 3.5 Normalization
3.6 Reward
Selecting an appropriate reward function is essential for a reinforcement algorithm to effectively learn the desired behavior, as emphasized by [69]. This section provides an overview of the reward functions proposed for the various L2RPN challenges.
@@ 3.6 Reward
3.7 Imitation learning
As shown in Table 1, eight out of twenty-two solutions employ Imitation Learning (IL), [15, 70, 71], a machine learning technique in which a task is learned by observing an expert. In IL, the agent tries to mimic the behavior of an expert by predicting the actions $a(t)\in A$ the expert would choose in state $s(t)\in S$. This differs from RL, where the agent learns through exploration, receiving rewards to guide it towards a policy that maximizes long-term cumulative reward. Although RL is effective in many scenarios, learning an optimal policy can be difficult, for example, when the rewards are sparse (e.g., given only at the end of an episode), or the reward function is undefined, which is the case in power grid maintenance. As discussed earlier, different reward functions have been designed to provide the agent with the right feedback. However, manually designing a reward function that satisfies the desired behavior can be extremely complicated. IL could be a good solution to address this issue. Moreover, even when a proper reward function is available, IL can speed up the learning process by reducing the number of samples required for RL.
IL has two main classes: behavioral cloning (BC) [72, 73] and inverse reinforcement learning (IRL) [74]. In BC, data trajectories $(\tau^{*})$ of the expert agent are collected, consisting of state-action pairs $(s_{0},a_{0}^{*}),(s_{1},a_{1}^{*}),\dots$, and supervised learning is used to learn the policy $\pi_{\theta}$, parameterized by $\theta$ using a neural network, that minimizes the loss $\mathcal{L}\left(a^{*},\pi_{\theta}(s)\right)$. In contrast, IRL tries to determine the reward function $r_{\theta}(s,a)$ that the expert agent is optimizing, based on observed demonstrations, rather than directly learning the policy.
@@ 3.7 Imitation learning
3.8 Prioritized replay
In its simplest form, RL algorithms discard experience after one update. Experience replay, a technique designed for off-policy algorithms, addresses two key issues: the risk of strongly correlated updates, and the tendency to forget useful but rare experiences immediately. Transitions, $(s_{t},a_{t},r_{t},s_{t+1})$, experienced by the agent, are stored in a replay buffer and sampled uniformly at random to be revisited. In prioritized experience replay (PER) [17], the idea is to replay transitions that are important more frequently and, in this way, make more efficient and effective use of the stored experiences. The priority is typically determined by the temporal difference (TD) error, which measures the difference between the predicted Q-value of a state-action pair and the target Q-value based on the current policy. Essentially, the TD error indicates how much the new experience impacts the learning process. As mentioned in [17], this prioritization might introduce a bias, which is corrected through importance sampling.
Several solutions in Table 1 that employ DQN-based algorithms, DDQN-2019, D3QN-2020 and D3QN-2022[12, 32, 50], make use of PER. In [50], the authors demonstrate the performance improvement when using PER for D3QN applied to the L2RPN problem.
The authors of PowRL-2022[52] also applied PER in their solution, but combined with PPO. This is noteworthy since PPO is an on-policy algorithm, and as previously mentioned, experience replay is primarily designed for off-policy methods. The use of PER in this context shows potential benefits outside its conventional use.
3.9 Exploration strategies
A DRL agent needs exploration to learn more about the environment and discover potentially good actions. A common exploration strategy used for off-policy algorithms is Epsilon-Greedy. Here, the agent selects a random action with probability $\epsilon$ and with probability $1-\epsilon$ the action best evaluated, $a^{*}=\operatorname*{arg\,max}_{a\in A}\{Q(s_{t},a)\}$, according to the current estimated Q-values.
Due to the large action space it can be difficult for the agent to converge to a good policy. DDQN-2019 and D3QN-2020[12, 32], therefore, introduce a guided exploration method. In this approach, the agent selects the top $N$ estimated Q-values for the current state $Q(s_{t},a)$ and simulates the performance of each action to pick the action with the highest reward. This leads to a more stable training process and significantly increases training efficiency. It should be noted that this approach uses a greedy action selection based on simulations and, therefore, might overlook actions that do not directly yield a high reward but do result in a new state $s_{t+1}$ for which the cumulative discounted future reward is profitable.
3.10 Curriculum learning or prioritized training
@@
A3C-2019[11] apply this technique by defining three difficulty levels and training the agent from easy to hard. The difficulty levels are defined by adjusting the Grid2Op [^22] The research of [11] has been applied on the predecessor of Grid2Op, PyPowNet. We discuss the equivalent parameters here. configuration parameters: SOFT_OVERFLOW_THRESHOLD ($SOT$), NB_TIMESTEP_OVERFLOW_ALLOWED ($TOA$ [^23] In [11] referred to as consecutive overload limit, COL.) and HARD_OVERFLOW_THRESHOLD ($HOT$). When a transmission line $e\in E$ has a soft overflow, i.e., $\rho_{e}(t)>SOT$, for more than $TOA$ consecutive time steps, the line will be disconnected. In case $\rho_{e}(t)>HOT$, the line $e$ immediately disconnects. In level-1 (the easiest) all parameters, $SOT$, $TOA$, and $HOT$ are set in such a way that lines would never disconnect, level-3 applies the default parameters of the environment, which are $SOT=1.0$, $TOA=3$ time steps and $HOT=1.5$ and level-2 is set in such a way that the strictness of the environment gradually increases. The authors show that adding this curriculum training strategy to the A3C algorithm applied to PNC enhances the training progress and improves the results of the final agent when tested on the evaluation episodes.
The authors of BDQN-2024[56] introduce K-means based predefined curriculum learning. The initial states that can be used as a start during the training process are grouped into clusters using K-means. One of the clusters is then sampled based on a calculated sampling weight, which is defined by the cluster’s difficulty and visiting frequency. With this sampling weight, the authors try to balance learning in a curriculum strategy, from easy to hard, while ensuring state variety.
Another solution that orders the data given to the DRL agent is SMAAC-2021[18]. In this approach, the training algorithm gives priority to more difficult episodes. Contrary to what curriculum learning suggests. The difficulty of an episode is determined by how well the do-nothing agent performed and how well the DRL agent has done so far. This has the advantage that the agent enters situations where it could learn more frequently, since difficult scenarios require more interference from the agent, whereas too easy scenarios might barely exceed the thermal limit of the transmission lines.
3.11 State and action space factorization
@@ 3.13 Activation threshold
3.14 Line reconnections and disconnections
@@
3.15 Revert to reference topology
@@
0.95 SAS-2021242424Based on code., D3QN-202024, PowRL-2022
0.9 AlphaZero-202224, LJNAgent-2024, CCMA-2025
0.85 BruteForce-202224
@@ 5.1 General setup
  • The action space is $\mathcal{A}_{n-0}$, described in Section 3.3,
@@ 5.2 Action space experiments
As mentioned in Section 3.3, most approaches reduce the action space to manage complexity in larger networks. The baseline Greedy and PPO agents use the $\mathcal{A}_{n-0}$ action space, which retains approximately 75% of the original $\mathcal{A}_{sym}$ action space (Table 3). A more restrictive reduction, $\mathcal{A}_{n-1}$, results in an action space of 40% of $\mathcal{A}_{sym}$. However, in the larger cases, 36 and 118, the action spaces are often reduced to around 1%–2%, most often using a Greedy reduction method, as described in Section 3.3. Therefore, in the experiments, we will also consider the smallest action space for case 14, containing roughly 5% of $\mathcal{A}_{sym}$, proposed by D3QN-2022, which we denote by $\mathcal{A}_{d3qn}$.
@@ 5.4 Reward function experiments
It is essential to choose a suitable reward function to help the RL algorithm learn the right behavior, as mentioned in Section 3.6. LABEL:Table:Rewards shows how each proposed solution uses a (sometimes slightly) different reward function. In the experiments, the following reward functions are compared for the training process of a PPO agent:
@@ 5.5 Curriculum training
Different from previous design choices, curriculum training is an optional enhancement of the RL algorithm. There are many possibilities for how to apply curriculum training to the PNC problem. We compare the baseline PPO agent with a curriculum-trained PPO agent using the method proposed by [11], described in Section 3.10. The configuration parameters for the three levels, $SOT$, $TOA$, $HOT$, discussed in Section 3.10, are shown in Table 12, along with an additional parameter $NOD$:NO_OVERFLOW_DISCONNECTION. With this setup, level 1 corresponds to a level where there is no line limit enforcement, level 2 disconnects after 15 time steps of high overload of 2, and level 3 corresponds to the default environment behavior [^28] The exact setup slightly deviates from [11] as the Grid2Op environment changed over time..
@@ 5.6 Rule-based experiments
As mentioned in Section 3.13, there is no golden rule to choose the activation threshold $\rho_{\textrm{act}}$. This motivates us to examine its impact by testing different values of $\rho_{\textrm{act}}$. We tested different values $\rho_{\textrm{act}}\in\{0.8,0.9,0.95,0.99\}$, evaluating both the training and validation performance in eight combinations of the training and final agent thresholds.
@@ 6.6 Activation threshold
We remark that when the threshold is higher, the agent interacts less with the environment per episode. Therefore, as illustrated in Fig. 10, training with a higher threshold also takes longer. Figs. 9 and 10, suggest that the training curve for $\rho_{\textrm{act}}=0.8$ may continue to improve in the no-opponent setting. To verify whether longer training improves performance, we conducted an additional experiment in which all agents were trained for the same duration, see Fig. 11. However, the upward trend for $\rho_{\textrm{act}}=0.8$ did not persist in the extended training experiment.
@@ 6.11 Evaluating the Rainbow Agent
From Figs. 14 and 15 we see that the Rainbow agent, configured with all best found configurations, indeed improves the mean episode survival time during training for both the PPO and the PPO agent, when compared to the Baseline agent. The comparable training curves of the ablation studies reveal that the performance of reverting to a single configuration closely matches that of the Rainbow agent. For example, we see that the impact of not including the change in activation threshold (Rainbow - $\rho_{\textrm{act}}=0.95$) is not as big as one might expect based on the previous results for the activation threshold.
@@ Appendix A Acronyms
Additinoally, Table 34 provides an overview of the other acronyms used throughout this paper.

### v2

@@ 3.1.3 Observation space
Active power Reactive power Voltages Cooldown Solution Acronym Complete Observation Time stamp Loads ($load\_p$) Gens ($gen\_p$) Lines ($p\_ex,p\_or$) Loads ($load\_qp$) Gens ($gen\_q$) Lines ($q\_ex,q\_or$) Loads ($load\_v$) Gens ($gen\_v$) Lines ($v\_ex,v\_or$) Line flows ($a\_{or},a\_{ex}$) Line loads ($\rho$) Time step overflow Line status Topo config ($topo\_vect$) line sub Maintenance Extra A3C-2019 DDQN-2019 A3C-2020191919Based on code, not (clearly) documented SMAAC-202119 Danger and history D3QN-202019 Binbinchen-2020 SAS-202119 CEM-2021 D3QN-2022 PowRL-2022 AlphaZero-2022 Curriculum-202319 HRL-2023 MARL-2023 Danger and history LJNAgent-202419 IL-2024 Thermal limit BDQN-2024 Zonal-2024 GNNIL-2025 Thermal limit CCMA-2025 Actions and action values.
@@ 3.1.4 Normalization
3.1.5 Reward
Selecting an appropriate reward function is essential for a reinforcement algorithm to effectively learn the desired behavior, as emphasized by [90]. This section provides an overview of the reward functions proposed for the various L2RPN challenges. As mentioned in Section 1.3, reward can be a function of the previous state, the action and the new state. All rewards discussed in this section are based only on the state $s_{t}$ observed at time $t$, and simply denoted by $r(t)$.
@@ 3.1.5 Reward
3.1.6 Imitation learning
As shown in Table 1, eight out of twenty-two solutions employ Imitation Learning (IL), see [42, 91, 92], a machine learning technique in which a task is learned by observing an expert. In IL, the agent tries to mimic the behavior of an expert by predicting the actions $a(t)\in A$ the expert would choose in state $s(t)\in S$. This differs from RL, where the agent learns through exploration, receiving rewards to guide it towards a policy that maximizes long-term cumulative reward. Although RL is effective in many scenarios, learning an optimal policy can be difficult, for example, when the rewards are sparse (e.g., given only at the end of an episode), or the reward function is undefined, which is the case in power grid maintenance. As discussed earlier, different reward functions have been designed to provide the agent with the right feedback. However, manually designing a reward function that satisfies the desired behavior can be extremely complicated. IL could be a good solution to address this issue. Moreover, even when a proper reward function is available, IL can speed up the learning process by reducing the number of samples required for RL.
IL has two main classes: behavioral cloning (BC), described in [93, 94], and inverse reinforcement learning (IRL), introduced in [95]. In BC, data trajectories $(\tau^{*})$ of the expert agent are collected, consisting of state-action pairs $(s_{0},a_{0}^{*}),(s_{1},a_{1}^{*}),\dots$, and supervised learning is used to learn the policy $\pi_{\theta}$, parameterized by $\theta$ using a neural network, that minimizes the loss $\mathcal{L}\left(a^{*},\pi_{\theta}(s)\right)$. In contrast, IRL tries to determine the reward function $r_{\theta}(s,a)$ that the expert agent is optimizing, based on observed demonstrations, rather than directly learning the policy.
@@ 3.1.6 Imitation learning
3.1.7 Prioritized replay
In its simplest form, RL algorithms discard experience after one update. Experience replay, a technique designed for off-policy algorithms (see Section 1.3.2), addresses two key issues: the risk of strongly correlated updates, and the tendency to forget useful but rare experiences immediately. Transitions, $(s_{t},a_{t},r_{t},s_{t+1})$, experienced by the agent, are stored in a replay buffer and sampled uniformly at random to be revisited. In prioritized experience replay (PER) [43], the idea is to replay transitions that are important more frequently and, in this way, make more efficient and effective use of the stored experiences. The priority is typically determined by the temporal difference (TD) error, which measures the difference between the predicted Q-value of a state-action pair and the target Q-value based on the current policy. Essentially, the TD error indicates how much the new experience impacts the learning process. As mentioned in [43], this prioritization might introduce a bias, which is corrected through importance sampling.
Several solutions in Table 1 that employ DQN-based algorithms, DDQN-2019, D3QN-2020 and D3QN-2022([39, 56, 71]), make use of PER. In [71], the authors demonstrate the performance improvement when using PER for D3QN applied to the L2RPN problem.
The authors of PowRL-2022([73]) also applied PER in their solution, but combined with PPO. This is noteworthy since PPO is an on-policy algorithm, and as previously mentioned, experience replay is primarily designed for off-policy methods. The use of PER in this context shows potential benefits outside its conventional use.
3.1.8 Exploration strategies
A DRL agent needs exploration to learn more about the environment and discover potentially good actions. A common exploration strategy used for off-policy algorithms (see Section 1.3.2) is Epsilon-Greedy. Here, the agent selects a random action with probability $\epsilon$ and with probability $1-\epsilon$ the action best evaluated, $a^{*}=\operatorname*{arg\,max}_{a\in A}\{Q(s_{t},a)\}$, according to the current estimated $Q$-values.
Due to the large action space it can be difficult for the agent to converge to a good policy. DDQN-2019 and D3QN-2020, therefore, introduce a guided exploration method. In this approach, the agent selects the top $N$ estimated $Q$-values for the current state $Q(s_{t},a)$ and simulates the performance of each action to pick the action with the highest reward. This leads to a more stable training process and significantly increases training efficiency. It should be noted that this approach uses a greedy action selection based on simulations and, therefore, might overlook actions that do not directly yield a high reward but do result in a new state $s_{t+1}$ for which the cumulative discounted future reward is profitable.
3.1.9 Curriculum learning or prioritized training
@@
A3C-2019([38]) apply this technique by defining three difficulty levels and training the agent from easy to hard. The difficulty levels are defined by adjusting the Grid2Op [^23] The research of [38] has been applied on the predecessor of Grid2Op, PyPowNet. We discuss the equivalent parameters here. configuration parameters: SOFT_OVERFLOW_THRESHOLD ($SOT$), NB_TIMESTEP_OVERFLOW_ALLOWED ($TOA$ [^24] In [38] referred to as consecutive overload limit, COL.) and HARD_OVERFLOW_THRESHOLD ($HOT$). When a transmission line $e\in E$ has a soft overflow, i.e., $\rho_{e}(t)>SOT$, for more than $TOA$ consecutive time steps, the line will be disconnected. In case $\rho_{e}(t)>HOT$, the line $e$ immediately disconnects. In level-1 (the easiest) all parameters, $SOT$, $TOA$, and $HOT$ are set in such a way that lines would never disconnect, level-3 applies the default parameters of the environment, which are $SOT=1.0$, $TOA=3$ time steps and $HOT=1.5$ and level-2 is set in such a way that the strictness of the environment gradually increases. The authors show that adding this curriculum training strategy to the A3C algorithm applied to PNC enhances the training progress and improves the results of the final agent when tested on the evaluation episodes.
The authors of BDQN-2024([77]) introduce K-means based predefined curriculum learning. The initial states that can be used as a start during the training process are grouped into clusters using K-means. One of the clusters is then sampled based on a calculated sampling weight, which is defined by the cluster’s difficulty and visiting frequency. With this sampling weight, the authors try to balance learning in a curriculum strategy, from easy to hard, while ensuring state variety.
Another solution that orders the data given to the DRL agent is SMAAC-2021([44]). In this approach, the training algorithm gives priority to more difficult episodes. Contrary to what curriculum learning suggests. The difficulty of an episode is determined by how well the do-nothing agent performed and how well the DRL agent has done so far. This has the advantage that the agent enters situations where it could learn more frequently, since difficult scenarios require more interference from the agent, whereas too easy scenarios might barely exceed the thermal limit of the transmission lines.
3.1.10 State and action space factorization
@@ 3.2.1 Activation threshold
3.2.2 Line reconnections and disconnections
@@
3.2.3 Revert to reference topology
@@
0.95 SAS-2021252525Based on code., D3QN-202025, PowRL-2022
0.9 AlphaZero-202225, LJNAgent-2024, CCMA-2025
0.85 BruteForce-202225
@@ 5.1 General setup
  • The action space is $\mathcal{A}_{n-0}$, described in Section 3.1.2,
@@ 5.2 Action space experiments
As mentioned in Section 3.1.2, most approaches reduce the action space to manage complexity in larger networks. The baseline Greedy and PPO agents use the $\mathcal{A}_{n-0}$ action space, which retains approximately 75% of the original $\mathcal{A}_{sym}$ action space (Table 3). A more restrictive reduction, $\mathcal{A}_{n-1}$, results in an action space of 40% of $\mathcal{A}_{sym}$. However, in the larger cases, 36 and 118, the action spaces are often reduced to around 1%–2%, most often using a Greedy reduction method, as described in Section 3.1.2. Therefore, in the experiments, we will also consider the smallest action space for case 14, containing roughly 5% of $\mathcal{A}_{sym}$, proposed by D3QN-2022, which we denote by $\mathcal{A}_{d3qn}$.
@@ 5.4 Reward function experiments
It is essential to choose a suitable reward function to help the RL algorithm learn the right behavior, as mentioned in Section 3.1.5. LABEL:Table:Rewards shows how each proposed solution uses a (sometimes slightly) different reward function. In the experiments, the following reward functions are compared for the training process of a PPO agent:
@@ 5.5 Curriculum training
Different from previous design choices, curriculum training is an optional enhancement of the RL algorithm. There are many possibilities for how to apply curriculum training to the PNC problem. We compare the baseline PPO agent with a curriculum-trained PPO agent using the method proposed by [38], described in Section 3.1.9. The configuration parameters for the three levels, $SOT$, $TOA$, $HOT$, discussed in Section 3.1.9, are shown in Table 12, along with an additional parameter $NOD$:NO_OVERFLOW_DISCONNECTION. With this setup, level 1 corresponds to a level where there is no line limit enforcement, level 2 disconnects after 15 time steps of high overload of 2, and level 3 corresponds to the default environment behavior [^29] The exact setup slightly deviates from [38] as the Grid2Op environment changed over time..
@@ 5.6 Rule-based experiments
As mentioned in Section 3.2.1, there is no golden rule to choose the activation threshold $\rho_{\textrm{act}}$. This motivates us to examine its impact by testing different values of $\rho_{\textrm{act}}$. We tested different values $\rho_{\textrm{act}}\in\{0.8,0.9,0.95,0.99\}$, evaluating both the training and validation performance in eight combinations of the training and final agent thresholds.
@@ 6.6 Activation threshold
We remark that when the threshold is higher, the agent interacts less with the environment per episode. Therefore, as illustrated in Fig. 13, training with a higher threshold also takes longer. Figs. 12 and 13, suggest that the training curve for $\rho_{\textrm{act}}=0.8$ may continue to improve in the no-opponent setting. To verify whether longer training improves performance, we conducted an additional experiment in which all agents were trained for the same duration, see Fig. 14. However, the upward trend for $\rho_{\textrm{act}}=0.8$ did not persist in the extended training experiment.
@@ 6.11 Evaluating the Rainbow Agent
From Figs. 17 and 18 we see that the Rainbow agent, configured with all best found configurations, indeed improves the mean episode survival time during training for both the PPO and the PPO agent, when compared to the Baseline agent. The comparable training curves of the ablation studies reveal that the performance of reverting to a single configuration closely matches that of the Rainbow agent. For example, we see that the impact of not including the change in activation threshold (Rainbow - $\rho_{\textrm{act}}=0.95$) is not as big as one might expect based on the previous results for the activation threshold.
@@ 8 Conclusion
Nevertheless, given the progress already achieved, the rapid growth of research in RL, and the many underexplored techniques with potential, such as imitation learning, off-policy techniques, curriculum learning, different neural network architectures and model-based RL mentioned in Section 7, we believe RL remains a promising direction to support grid operators in PNC.
@@ Appendix A Acronyms
In addition, Table 34 provides an overview of the other acronyms used throughout this paper.